OpenStudy (anonymous):

Determine if the given lines are parallel, perpendicular or neither: 3x + 4y = 6 and 4y = 3x + 7.

OpenStudy (anonymous):

well if they're parallel they will have the same slope, and a different Y-intercept we can see the y intercept are different, but what about the slopes? are they the same?

OpenStudy (anonymous):

Oh My God... My head almost exploded and everything just went blank in my mind. I have two equations. Is the first step supposed to be to determine what the X and Y

OpenStudy (anonymous):

wait, no need to panic... the fist step is to determine the slope ...rise over run.. right

OpenStudy (anonymous):

no, just put both equations in slope intercept form

OpenStudy (anonymous):

what does that mean again

OpenStudy (anonymous):

Y= 3x-6/3

OpenStudy (anonymous):

Y = mx + b

OpenStudy (anonymous):

3x + 4y = 6 subtract 3x from each side 4y = -3x + 6 then divide each side by 4 then it is in slope intercept form :]

OpenStudy (anonymous):

I got: Y= -3/4 + 1 1/2

OpenStudy (anonymous):

y = -3/4x + 6/4 6/4 simplified would be 3/2

OpenStudy (anonymous):

what i did is to plot them using transformation of equation: for the first one i plotted x and y = (6-3x)/4 and for the second one i plotted x and y = (3x+7)/4....they are neither parallel or perpendicular...they looked like a cross or multiplication symbol on the chart like this : "x"

OpenStudy (anonymous):

and Y = -3x/4 + 7/4

OpenStudy (anonymous):

Therefore, the lines neither cross nor run parallel

OpenStudy (anonymous):

they do cross...they become like this : "X"

OpenStudy (anonymous):

y = 3/4x + 7/4 the slope is positive that's why we get the X due to the slopes being different :]

OpenStudy (anonymous):

they would be perpendicular if they cross at 90 deg but they do not!

OpenStudy (anonymous):

so the first one (3x + 4y=6) is a negative slope

OpenStudy (anonymous):

@ChristaB yes because we're subtracting 3x from each side that's why it becomes negative

OpenStudy (anonymous):

@vick2075 it would be if they had the same y-intercept as well, i made an error, and saying they cross at 90deg makes no sense, do you mean form right angles? o.o

OpenStudy (anonymous):

yeah form right angles.....90 deg on both sides of the line with respect to the other line

OpenStudy (anonymous):

@vick2075 no...

OpenStudy (anonymous):

no...what?! anyway how do you put "@vick2075" at the front?

OpenStudy (anonymous):

@vick2075 hit the @ symbol then type some letter and hit the name you want

OpenStudy (anonymous):

@gaara438125 ....right! thx!

OpenStudy (anonymous):

@ChristaB have you got the chart?

OpenStudy (anonymous):

This has been working out amazing! Thank you guys SOOOo much. I'm ready for the next problem. So, I will close this one out and write out the next one. If you are up for it, I ask you to hang with me cause you are truly my hero's right now

OpenStudy (anonymous):

which chart

OpenStudy (anonymous):

@ChristaB the one i attached to one of my replies!

OpenStudy (anonymous):

hmm.. no. I didnt see a chart. how would I find that? Im new to this sight and Thank God I found it!!!!)

OpenStudy (anonymous):

@ChristaB it's on my second reply above

OpenStudy (anonymous):

Ok, that is a totally cool thing... (the attachment thing)

OpenStudy (anonymous):

It would not let me open the doc... but the fact that you can send an attachment is pretty nifty